# SemEval Task 1: Semantic Textual Relatedness for African and Asian Languages

### v1

@@ 1 Introduction
Prior shared tasks Agirre et al. (2012, 2013, 2014, 2015, 2016); Cer et al. (2017a) have mainly focused on textual similarity. In this work, we provide participants with SemRel Ousidhoum et al. (2024), a collection of 14 newly curated monolingual STR datasets for Afrikaans (afr), Amharic (amh), Modern Standard Arabic (arb), Algerian Arabic (arq), Moroccan Arabic (ary), English (eng), Spanish (esp), Hausa (hau), Hindi (hin), Indonesian (ind), Kinyarwanda (kin), Marathi (mar), Punjabi (pun) and Telugu (tel). The datasets are composed of sentence pairs, each assigned a relatedness score between 0 (completely unrelated) and 1 (maximally related) with a large range of expected relatedness values. The pairs of sentences were first selected from pre-existing datasets covering various topics and formality levels, e.g., news data, Wikipedia, and conversational data. To generate the relatedness scores, the sentence pairs were then annotated by native speakers who performed comparisons between different pairs of sentences using Best–Worst Scaling (BWS) Louviere and Woodworth (1991); Kiritchenko and Mohammad (2017a). The shared task included three main tracks: (1) supervised, (2) unsupervised, and (3) cross-lingual.
@@ 2 Related Work
Most prior shared tasks focus on semantic textual similarity, a narrower subset of relatedness, and often only cover high-resource languages such as English Agirre et al. (2012, 2013, 2014, 2015, 2016), Arabic, German, Spanish, and Turkish Cer et al. (2017b) with few exceptions such as Armendariz et al. (2020) who also included Slovene, Finnish, and Croatian.
@@ 4 Task Description
| Track A (supervised) | Track B (unsupervised) | Sub-task C (Crosslingual) | ||||
| # | Team | Score | Team | Score | Team | Score |
| * | Lexical Overlap | 0.456 | ||||
| * | SemRel baseline (LaBSE) | 0.762 | SemRel baseline (XLMR) | 0.353 | SemRel baseline (LaBSE) | 0.579 |
| 1 | AAdam | 0.800 | SATLab | 0.543 | AAdaM | 0.650 |
| 2 | NRK | 0.781 | MasonTigers | 0.514 | UAlberta | 0.589 |
| 3 | PEAR | 0.758 | HW–TSC | 0.482 | silp_nlp | 0.566 |
| 4 | silp_nlp | 0.740 | UAlberta | 0.481 | USTC_NLP | 0.523 |
| 5 | NLP_1@SSN | 0.740 | silp_nlp | 0.400 | MaiNLP | 0.499 |
@@ 5.4 Task A: Supervised
| 7 | MBZUAI-UNAM Barajas et al. (2024) | 0.840 | 0.541 | 0.786 | 0.832 | 0.697 | 0.670 | 0.458 | 0.867 | 0.785 | 0.720 |
@@
| NR | UMBCLU Dipta and Vallurupalli (2024) | - | - | 0.745 | 0.838 | 0.721 | 0.640 | 0.681 | 0.841 | 0.682 | 0.733 |
@@
| NR | Pinealai Eponon and Luis Perez (2024) | - | - | - | 0.837 | - | - | - | - | - | 0.837 |
@@ 5.6 Task C: Crosslingual
| 4 | USTC_NLP Li et al. (2024a) | 0.749 | 0.709 | 0.517 | 0.414 | 0.613 | 0.784 | 0.685 | 0.476 | 0.658 | 0.460 | 0.454 | -0.248 | 0.523 |
| 5 | MaiNLP Zhou et al. (2024) | 0.738 | 0.728 | 0.399 | 0.274 | 0.568 | - | - | - | 0.695 | 0.319 | 0.681 | 0.087 | 0.499 |
| 6 | ustcctsu Li et al. (2024a) | 0.603 | 0.656 | 0.469 | 0.420 | 0.402 | 0.700 | 0.689 | 0.111 | 0.596 | 0.476 | 0.302 | -0.084 | 0.445 |
| 7 | umbclu Dipta and Vallurupalli (2024) | 0.822 | 0.043 | 0.035 | 0.126 | -0.038 | 0.788 | 0.609 | 0.457 | 0.155 | 0.515 | 0.484 | -0.078 | 0.326 |
| 8 | HausaNLP Salahudeen et al. (2024) | 0.737 | -0.031 | 0.184 | 0.074 | 0.276 | 0.360 | 0.604 | 0.177 | 0.346 | 0.472 | 0.319 | 0.114 | 0.303 |
| 9 | MasonTigers Goswami et al. (2024) | 0.385 | 0.131 | 0.213 | 0.221 | 0.203 | 0.310 | 0.557 | 0.099 | 0.511 | 0.133 | 0.079 | 0.020 | 0.239 |
@@ 5.6.1 Best Performing Systems
They used an XGBoost regressor-based (Chen and Guestrin, 2016) ensemble approach to integrate the predicted relatedness scores of three distinct regression models, with one optional SBERT model, as input and returned the final relatedness score as output. They applied the English version of their method trained for Track A to the translations of the non-English test sets. The regression model fine-tuned on MPNet was used in the XGBoost ensemble only for amh, hau, and hin, but not for the others, including esp, ary, kin, ind, arb, arq, and afr. The pre-trained English language models that were used include RoBERTa Large, T5 Base, and GPT2 Base, as well as MPNet only for languages amh, hau, and hin.
@@ Appendix A Appendix: Track A–Best Performing Teams
They used an ensemble approach with an XGBoost regression (Chen and Guestrin, 2016) to integrate the predicted relatedness scores of three distinct regression models, with one optional SBERT model, as inputs and to return the final relatedness scores as output. Each of these models used a different pre-trained language model as its backbone, specifically RoBERTa Large (Liu et al., 2019b), T5 Base, GPT-2 Base, and the optional SBERT (MPNet).
They also translated the English data to esp, hau, amh, mar, tel and used it for each language, respectively. They merged the dev and training sets into one set, then split them again via uniform random sampling according to their original sizes to establish new training and development splits. The did not use the data provided for arq, ary, and kin, and applied the English-trained version of their method to the English translations of the arq, ary, and kin test sets instead.
@@ Appendix C Appendix: Track C
They applied the English version of their method reported for Track A to the translations of the non-English test sets. The regression model fine-tuned on MPNet was used in the XGBoost ensembling method for amh, hau, and hin. However, for esp, ary, kin, ind, arb, arq, and afr. For Track C, pre-trained English language models were used, including RoBERTa Large, T5 Base, and GPT2 Base, as well as MPNet for amh, hau, hin.
@@
| MBZUAI–UNAM | Barajas et al. (2024) |
@@
| Pinealai | Eponon and Luis Perez (2024) |
@@
| UMBCLU | Dipta and Vallurupalli (2024) |

### v5

# SemEval-2024 Task 1: Semantic Textual Relatedness for African and Asian Languages

@@ 1 Introduction
Prior shared tasks Agirre et al. (2012, 2013, 2014, 2015, 2016); Cer et al. (2017) have mainly focused on textual similarity. In this work, we provide participants with SemRel Ousidhoum et al. (2024), a collection of 14 newly curated monolingual STR datasets for Afrikaans (afr), Amharic (amh), Modern Standard Arabic (arb), Algerian Arabic (arq), Moroccan Arabic (ary), English (eng), Spanish (esp), Hausa (hau), Hindi (hin), Indonesian (ind), Kinyarwanda (kin), Marathi (mar), Punjabi (pun) and Telugu (tel). The datasets are composed of sentence pairs, each assigned a relatedness score between 0 (completely unrelated) and 1 (maximally related) with a large range of expected relatedness values. The pairs of sentences were first selected from pre-existing datasets covering various topics and formality levels, e.g., news data, Wikipedia, and conversational data. To generate the relatedness scores, the sentence pairs were then annotated by native speakers who performed comparisons between different pairs of sentences using Best–Worst Scaling (BWS) Louviere and Woodworth (1991); Kiritchenko and Mohammad (2017a). The shared task included three main tracks: (1) supervised, (2) unsupervised, and (3) cross-lingual.
@@ 2 Related Work
Most prior shared tasks focus on semantic textual similarity, a narrower subset of relatedness, and often only cover high-resource languages such as English Agirre et al. (2012, 2013, 2014, 2015, 2016), Arabic, German, Spanish, and Turkish Cer et al. (2017) with few exceptions such as Armendariz et al. (2020) who also included Slovene, Finnish, and Croatian.
@@ 4 Task Description
| Track A (Supervised) | Track B (Unsupervised) | Track C (Crosslingual) | ||||||
| # | Team | Score | Team | Score | Team | Score | ||
| * | Lexical Overlap | 0.456 | ||||||
| * | baseline (LaBSE) | 0.762 | * | baseline (XLMR) | 0.353 | * | baseline (LaBSE) | 0.579 |
| 1 | AAdam | 0.800 | SATLab | 0.543 | AAdaM | 0.650 | ||
| 2 | NRK | 0.781 | MasonTigers | 0.514 | UAlberta | 0.589 | ||
| 3 | PEAR | 0.758 | HW–TSC | 0.482 | silp_nlp | 0.566 | ||
| 4 | silp_nlp | 0.740 | UAlberta | 0.481 | MaiNLP | 0.499 | ||
| 5 | NLP_1@SSN | 0.740 | silp_nlp | 0.400 | ustcctsu | 0.445 | ||
@@ 5.4 Task A: Supervised
| 7 | MBZUAI-UNAM Ortiz-Barajas et al. (2024) | 0.840 | 0.541 | 0.786 | 0.832 | 0.697 | 0.670 | 0.458 | 0.867 | 0.785 | 0.720 |
@@
| NR | UMBCLU Roy Dipta and Vallurupalli (2024) | - | - | 0.745 | 0.838 | 0.721 | 0.640 | 0.681 | 0.841 | 0.682 | 0.733 |
@@
| NR | Pinealai Eponon and Ramos Perez (2024) | - | - | - | 0.837 | - | - | - | - | - | 0.837 |
@@ 5.6 Task C: Crosslingual
| 4 | MaiNLP Zhou et al. (2024) | 0.738 | 0.728 | 0.399 | 0.274 | 0.568 | - | - | - | 0.695 | 0.319 | 0.681 | 0.087 | 0.499 |
| 5 | USTCCTSU Li et al. (2024a) | 0.603 | 0.656 | 0.469 | 0.420 | 0.402 | 0.700 | 0.689 | 0.111 | 0.596 | 0.476 | 0.302 | -0.084 | 0.445 |
| 6 | umbclu Roy Dipta and Vallurupalli (2024) | 0.822 | 0.043 | 0.035 | 0.126 | -0.038 | 0.788 | 0.609 | 0.457 | 0.155 | 0.515 | 0.484 | -0.078 | 0.326 |
| 7 | HausaNLP Salahudeen et al. (2024) | 0.737 | -0.031 | 0.184 | 0.074 | 0.276 | 0.360 | 0.604 | 0.177 | 0.346 | 0.472 | 0.319 | 0.114 | 0.303 |
| 8 | MasonTigers Goswami et al. (2024) | 0.385 | 0.131 | 0.213 | 0.221 | 0.203 | 0.310 | 0.557 | 0.099 | 0.511 | 0.133 | 0.079 | 0.020 | 0.239 |
| NR | USTC_NLP | 0.749 | 0.709 | 0.517 | 0.414 | 0.613 | 0.784 | 0.685 | 0.476 | 0.658 | 0.460 | 0.454 | -0.248 | 0.523 |
@@ 5.6.1 Best Performing Systems
They used an XGBoost regressor-based (Chen and Guestrin, 2016) ensemble approach to integrate the predicted relatedness scores of three distinct regression models, with one optional SBERT model, as input and returned the final relatedness score as output. They applied the English version of their method trained for Track A to the translations of the non-English test sets. The regression model fine-tuned on MPNet was used in the XGBoost ensemble only for amh, hau, and hin, but not for the other languages such as esp, ary, kin, ind, arb, arq, and afr. The pre-trained English language models that were used include RoBERTa Large, T5 Base, and GPT2 Base, as well as MPNet only for languages amh, hau, and hin.
@@ Appendix A Appendix: Track A–Best Performing Teams
They used an ensemble approach with an XGBoost regression (Chen and Guestrin, 2016) to integrate the predicted relatedness scores of three distinct regression models, with one optional SBERT model, as input and returned the final relatedness scores as output. Each of these models used a different pre-trained language model as its backbone, specifically RoBERTa Large (Liu et al., 2019b), T5 Base, GPT-2 Base, and the optional SBERT (MPNet). They merged the English training and development sets with the translated training set of the target language. Then, they split them again via uniform random sampling according to their original sizes to establish new training and development splits. The did not use the data provided for arq, ary, and kin, and applied the English-trained version of their method to the English translations of the arq, ary, and kin test sets instead.
@@ Appendix C Appendix: Track C
They applied the English version of their method reported for Track A to the translations of the non-English test sets. The regression model fine-tuned on MPNet was used in the XGBoost ensembling method for amh, hau, and hin and not for esp, ary, kin, ind, arb, arq, and afr.
@@
| MBZUAI–UNAM | Ortiz-Barajas et al. (2024) |
@@
| Pinealai | Eponon and Ramos Perez (2024) |
@@
| UMBCLU | Roy Dipta and Vallurupalli (2024) |